Given two complex numbers z and w such that z=(1+i)w+(3-i)w^_ where w^_is the conjugate of w which of the following is w^_ in terms of z and z^_


1
Expert's answer
2021-10-28T07:38:38-0400

"\displaystyle
z = (1+i)w+(3-i)\overline{w}\\
\text{Making $\overline{w}$ the subject of the formula, we have}\\
\overline{w} = \frac{z-(1+i)w}{3-i}\\
\text{Next we rationalize the fraction by multiplying the numerator and denominator by }\\
\text{3+i to obtain }\\
\frac{z(3+i)-(2+4i)}{10}"